5. Given the probability density f(x)= for -0<x<00, find k. 1+ 2 Jor -
Answer #1

The entire area of f(x) should be 1. Hence,

I sp + -

\small k*\int_{-\infty}^{\infty} \frac{1}{1+x^2}~dx= 1

k*(\small \pi) = 1

k = 1/\small \pi